### v1

# Enhancement of temperature of quantum anomalous Hall effect in two-dimensional germanene/magnetic semiconductor heterostructures

Abstract
Quantum anomalous Hall effect (QAHE) is significant for future low-power electronics devices, where a main challenge is realizing QAHE at high temperatures. In this work, based on experimentally reported two-dimensional (2D) germanene and magnetic semiconductors Cr2Ge2Te6 and Cr2Si2Te6, and the first principle calculations, germanene/magnetic semiconductor heterostructures are investigated. Topologically nontrivial edge states and quantized anomalous Hall conductance are demonstrated. It is shown that the QAHE temperature can be enhanced to approximately 62 K in germanene/monolayer (ML) Cr2Ge2Te6 with 2.1% tensile strain, 64 K in germanene/bilayer (BL) Cr2Ge2Te6 with 1.4% tensile strain, and 50 K in germanene/ML Cr2Si2Te6 with 1.3% tensile strain. With increasing tensile strain of these heterostructures, the band gap decreases and the Curie temperature rises, and the highest temperature of QAHE is obtained. Since these 2D materials were discovered in recent experiments, our results provide promising materials for achieving high-temperature QAHE.
I Introduction
Quantum anomalous Hall effect (QAHE) is one of the most exciting phenomenon of matter and is characterized by an integer Chern number $\mathcal{C}$ [1, 2]. A key manifestation of QAHE is the dissipationless chiral edge current. It has the potential to revolutionize future low-power quantum electronic devices. Therefore, much theoretical and experimental research has been conducted to achieve QAHE at high temperatures [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Various methods are being explored to achieve QAHE. Introducing ferromagnetic order in topological insulators (TIs) is a primary method for achieving QAHE. The first observation of QAHE occurred in magnetically doped TIs, i.e., Cr-doped (Bi,Sb)2Te3 thin films at 30 mK [3]. Subsequent research has focused on increasing the QAHE temperature. For example, the anomalous Hall resistance $\rho_{xy}$ reaches a maximum value of $0.95h/e^{2}$ in a 10 quintuple-layer (QL) thin film of Cr0.1(Bi0.5,Sb0.5)1.9Te3 at 280 mK [8], and a six QL (Cr0.12Bi0.26Sb0.62)2Te3 film exhibits a quantized anomalous Hall resistance $R_{yx}$ up to 0.3 K [10]. To date, the highest reported record QAHE temperature in these doped magnetic TIs is approximately 2 K [11]. Because disordered magnetic impurities can inevitably degrade sample quality, thus limiting the QAHE temperature. Many researchers are investigating intrinsic QAHE materials. In recent years, experimental breakthroughs in QAHE have been achieved in the van der Waals (vdW) layered material MnBi2Te4 [12, 13, 14]. MnBi2Te4 is a layered magnetic TI consisting of Te-Bi-Te-Mn-Te-Bi-Te septuple layers (SLs) with antiferromagnetic ground state characterized by antiferromagnetic interlayer coupling and ferromagnetic intralayer coupling. In five SL flakes, the quantization temperature can be increased to 6.5 K by aligning all layers ferromagnetically with an external magnetic field [12].
Another approach to achieving the QAHE state involves fabricating TI/magnetic semiconductor heterostructures. Recent experimental evidnece has shown that low-buckled epitaxial germanene is a quantum spin Hall insulator with a large bulk gap and robust metallic edges at 77 K [21]. The quantized spin Hall conductance has also been confirmed by first principle calculations [22]. Furthermore, two-dimensional (2D) Cr2Ge2Te6 has been experimentally demonstrated as a ferromagnetic semiconductor, with a Curie temperature ($T_{C}$) of 28 K for bilayer (BL) Cr2Ge2Te6 [23]. Previous theoretical studies suggest that chiral edge states and quantized Hall conductance can be realized in germanene/Cr2Ge2Te6 heterostructures [24, 25]. A key question is whether the QAHE temperature can be enhanced in germanene/monolayer (ML) Cr2Ge2Te6. Recent studies have indicated that tensile strain is an effective method for increasing the $T_{C}$ of 2D Cr2Ge2Te6 [26, 27]. First-principles calculations predict a $T_{C}$ of 144 K for ML Cr2Ge2Te6 under 5% tensile strain [27]. Experimentally, strain can be applied to samples through various techniques [28, 29, 30, 27, 31, 32], including using a micro-manipulator on the substrate to transfer strain to the sample [28] or exploiting lattice mismatch between different materials [29]. Gradient-modulated strain up to 4% has been reported in experiments with vdW heterostructures [32].
In this paper, we report our first-principles study on enhancing the QAHE temperature in germanene/magnetic semiconductor heterostructures. We investigated the effect of biaxial tensile strain on these heterostructures to increase the QAHE temperature. Our calculations show that a QAHE temperature of 62 K is achieved in the germanene/ML Cr2Ge2Te6 heterostructure with 2.1% tensile strain. For the germanene/BL Cr2Ge2Te6 heterostructure, a QAHE temperature of 64 K is obtained with 1.4% tensile strain, while the germanene/ML Cr2Si2Te6 heterostructure exhibits a QAHE temperature of 50 K under 1.3% tensile strain. These results suggest a promising route for achieving high-temperature QAHE using experimentally available 2D materials.
Figure 1: Crystal structures of two-dimensional germanene/ML Cr2Ge2Te6 heterostructure: (a) side view, and (b) top view, respectively. The arrows indicate biaxial tensile strain applied to the heterostructure.
II Method
[…]
III QAHE in germanene/ML Cr2Ge2Te6 heterostructure
The crystal structure with the side and top views of the germanene/ML Cr2Ge2Te6 heterostructure is shown in Fig. 1. The heterostructure is constructed using a $\sqrt{3}\times\sqrt{3}$ supercell of germanene and a $1\times 1$ unit cell of ML Cr2Ge2Te6 with an interlayer distance of $d=3.45$. The calculated lattice mismatch between germanene and Cr2Ge2Te6 is less than 1%. The optimized in-plane lattice constant of the germanene/Cr2Ge2Te6 heterostructure is $a=6.954$, which is slightly smaller than that of the optimized ML Cr2Ge2Te6 ($a=6.964$).
Figure 2: (a) Electronic band structures and (b) density of states of germanene/ML Cr2Ge2Te6 heterostructure obtained by the first principle calculations with SOC.
Figure 3: For germanene/ML Cr2Ge2Te6 heterostructure, (a) anomalous Hall conductivity and (b) chiral edge states near the Fermi level for the case without strain. (c)-(d) for the case with 2% tensile strain.
Figure 4: (a) Four different spin configurations for Cr atoms in the honeycomb lattice: FM (ferromagnetic), Neel AFM (antiferromagnetic), Stripy AFM and Zigzag AFM. Red and green arrows represent the Cr atom with spin up and spin down polarizations, respectively. (b) The calculated magnetization and susceptibility as a function of temperature for of germanene/Cr2Ge2Te6 heterostructure without applying strain by MC simulations, where the Curie temperature is obtained as $T^{MC}_{C}$ = 57 K. (c) The rescaled Curie temperature $T_{C}$ as a function of tensile strain by MC simulations for germanene/ML Cr2Ge2Te6 heterostructure and ML Cr2Ge2Te6. See text for details.
The electronic band structures and density of states (DOS) of germanene/ML Cr2Ge2Te6 heterostructure are shown in Fig. 2. With SOC included, the germanene/ML Cr2Ge2Te6 heterostructure exhibits a finite band gap of 46 meV. This indicates that even at room temperature, thermal fluctuations are insufficient to fully populate the band gap. The Dirac cone around the $\Gamma$ point is predominantly contributed by $p_{z}$ orbits of germanene. Notably, this band gap is larger than that of pristine germanene. Since SOC depends on the atomic number $(\sim Z^{4})$ and SOC strength of Te is larger than that of germanene, the observed band gap enhancement is attributed to interlayer coupling and hybridization. To investigate the topological properties, maximally localized Wannier functions (MLWFs) implemented in the Wannier90 code were used to fit our density functional theory (DFT) band calculations. The calculated Chern number for the valence band near the Fermi level is $C=1$, confirming the topologically nontrivial nature of the band structure. Fig. 3(a) shows the calculated anomalous Hall conductance as a function of chemical potential, revealing a quantized Hall plateau at $\sigma_{xy}=\frac{e^{2}}{h}$. The plateau is consistent with the presence of one topologically nontrivial edge state connecting the bulk states in Fig. 3(b).
To investigate magnetic properties of this heterostructure, we considered the following Heisenberg-type Hamiltonian:
$\begin{split}H=&\sum_{<i,j>}J_{1}\textbf{S}_{i}\cdot\textbf{S}_{j}+\sum_{<<i,j>>}J_{2}\textbf{S}_{i}\cdot\textbf{S}_{j}\\ &+\sum_{<<<i,j>>>}J_{3}\textbf{S}_{i}\cdot\textbf{S}_{j}+M\sum_{i}(S^{z}_{i})^{2},\end{split}$ (1)
where $J_{1}$, $J_{2}$, $J_{3}$ denote the nearest neighbor, second-nearest-neighbor and third-nearest-neighbor exchange couplings, respectively. To determine the magnetization direction, we preformed total-energy calculations for both out-of plane and in-plane magnetization. Our results show that the easy axis is perpendicular to the 2D plane, consistent with experimental results for 2D Cr2Ge2Te6 [23]. To obtain exchange couplings, we considered four spin configurations: ferromagnetic (FM), Neel antiferromagnetic (AFM), Stripy AFM and Zigzag AFM as shown in Fig. 4(a). The energy of four configurations are expressed as
$\begin{split}E_{FM}&=12J_{1}+24J_{2}+12J_{3}+E_{0},\\ E_{NAFM}&=-12J_{1}+24J_{2}-12J_{3}+E_{0},\\ E_{SAFM}&=-4J_{1}-8J_{2}+12J_{3}+E_{0},\\ E_{ZAFM}&=4J_{1}-8J_{2}-12J_{3}+E_{0},\end{split}$ (2)
where $E_{0}$ is the energy which is independent of the spin configurations. $J_{1}$, $J_{2}$, $J_{3}$ are provided in supplementary materials (SMs) [40]. The Monte Carlo (MC) simulations were performed on a 20 $\times$ 20 $\times$ 1 hexagonal lattice with periodic boundary conditions, using $10^{6}$ MC steps per temperature. The resulting temperature dependence of the magnetization and susceptibility for unstrained germanene/ML Cr2Ge2Te6 heterostructure is presented in Fig. 4(b).
To investigate the effect of tensile strain on the critical temperature of germanene/Cr2Ge2Te6 heterostructure, biaxial tensile strain was applied. Strain is defined as $(a-a_{0})/a_{0}$, where $a$ and $a_{0}$ are the lattice constants with and without strain, respectively. As shown in Fig. 4(c), $T_{C}$ increases with strain. To address the overestimation of $T_{C}$ by MC results, a rescaling method was employed. The calculated $T^{MC}_{C}$ of 50 K for ML Cr2Ge2Te6 without stain was scaled to match the experimental $T_{C}$ of 28 K for 2D Cr2Ge2Te6 [23]. All $T^{MC}_{C}$ values for strained germanene/ML Cr2Ge2Te6 and ML Cr2Ge2Te6 were then normalized by the unstrained ML Cr2Ge2Te6 $T^{MC}_{C}$ of 50 K. These ratios were subsequently multiplied by 28 K to provide more reliable $T_{C}$ estimates as shown in Fig. 4(c).
As shown in Fig. 5, increasing tensile strain in germanene/ML Cr2Ge2Te6 heterostructure leads to an increase in $T_{C}$ and a decrease in the band gap. The gap nearly closes at a tensile strain of 2.4%. The band gap have been given in unit of K. The yellow region in Fig. 5 indicates the predicted QAHE phase. Notably, a QAHE state is predicted up to 62 K in the germanene/ML Cr2Ge2Te6 heterostructure with 2.1% tensile strain.
Figure 5: The band gap and Curie temperature $T_{C}$ as a function of tensile strain of germanene/ML Cr2Ge2Te6 heterostructure.
To understand the decrease in band gap with increasing tensile strain, we analyzed the fatband structure around the Fermi level for germanene/ML Cr2Ge2Te6 heterostructure at strain=0% and strain=2% as shown in Fig. 6. Clearly, the bands around $\Gamma$ point are mainly contributed by $p_{z}$ orbitals of germanene, while those along the $\Gamma-K$ path are dominated by $d_{xz}$ and $d_{yz}$ orbitals of Cr. In the unstrained heterostructure, the direct band gap is located at $\Gamma$. With increasing tensile strain, the bands at $\Gamma$ show little variation, but the bands along $\Gamma-K$ shift downwards, resulting in a reduced band gap.
Figure 6: The fatband around the Fermi level of germanene/ML Cr2Ge2Te6 heterostructure (a) without strain and (b) with 2% tensile strain.
The great enhancement of $T_{C}$ in germanene/ML Cr2Ge2Te6 heterostructure under tensile strain can be explained by considering the superexchange interaction [41, 42, 43]. The superexchange interaction involves two processes. One is the direct exchange process between the $d_{2}$ electrons of the Cr2 atom and the $p^{\prime}$ electrons of the Te atom, presented by $J_{p^{\prime}d_{2}}$. The other is the hopping process between the $p$ electrons of the Te atom and the $d_{1}$ electrons of the Cr1 atom, presented by $V_{pd_{1}}$. The coupling $J_{super}$ between the two Cr atoms can be expressed as [40]
$\begin{split}J_{super}&=\frac{1}{4}\sum_{d_{1},p,p^{\prime},d_{2}}|V_{pd_{1}}|^{2}J_{p^{\prime}d_{2}}[\frac{1}{(E_{d_{1}d_{1}^{\prime}}^{\uparrow\uparrow})^{2}}-\frac{1}{(E_{d_{1}d_{1}^{\prime}}^{\uparrow\downarrow})^{2}}].\\ &=\frac{1}{4A}\sum_{d_{1},p,p^{\prime},d_{2}}|V_{pd_{1}}|^{2}J_{p^{\prime}d_{2}},\end{split}$ (3)
where $A\equiv 1/(1/(E_{d_{1}d_{1}^{\prime}}^{\uparrow\uparrow})^{2}-1/(E_{d_{1}d_{1}^{\prime}}^{\uparrow\downarrow})^{2})$ is taken as a pending parameter and does not change with strain. $E_{d_{1}d_{1}^{\prime}}^{\uparrow\uparrow}$ and $E_{d_{1}d_{1}^{\prime}}^{\uparrow\downarrow}$ are energies of two $d$ electrons at the same Cr atom with parallel and antiparallel spins, respectively. $V_{pd_{1}}$ is the hopping parameter between the $p$ electrons and the $d_{1}$ electrons. The direct p-d exchange $J_{p^{\prime}d_{2}}$ can be derived from the s-d exchange model following the Schrieffer-Wolff transformation [44]
$\begin{split}J_{p^{\prime}d_{2}}=2|V_{p^{\prime}d_{2}}|^{2}(\frac{1}{E_{p^{\prime}}-E_{d_{2}}}+\frac{1}{E_{d_{2}}+U-E_{p^{\prime}}}),\end{split}$ (4)
where $E_{p^{\prime}}$ and $E_{d_{2}}$ are the energy level of the $p^{\prime}$ electrons and the $d_{2}$ electrons, respectively. $V_{p^{\prime}d_{2}}$ is the hopping parameter between the $p^{\prime}$ electrons and the $d_{2}$ electrons. These parameters can be obtained by DFT calculation and Wannier90 code.
Due to the crystal field, the $d$ electron configuration of Cr is split into three groups: $a_{1}$ ($d_{z^{2}}$ orbital), $e_{1}$ ($d_{xy}$ and $d_{x^{2}-y^{2}}$ orbitals) and $e_{2}$ ($d_{xz}$ and $d_{yz}$ orbitals). Wannier90 calculations indicate that the main contribution comes from $e_{2}$. The values of $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ for the Cr1-Te1-Cr2 path in ML Cr2Ge2Te6 without and with tensile strain are listed in Table. 1. The resulting superexchange interaction $J_{super}$ for relaxed and strained ML Cr2Ge2Te6 are $0.372/A$ and $0.878/A$, respectively. The ratio of $J_{super}$ for strained to relaxed ML Cr2Ge2Te6 is 2.36, which is comparable to the ratio (1.25) of the DFT-calculated nearest neighbor exchange coupling $J_{1}$ for the strained and relaxed germanene/ML Cr2Ge2Te6 heterostructure. Comparison of the strained and unstrained ML Cr2Ge2Te6 results reveals that the enhancement of the superexchange interaction $J_{super}$ is primarily due to the decreased energy difference $E_{p^{\prime}}-E_{d_{2}}$. This decrease in $E_{p^{\prime}}-E_{d_{2}}$ contributes to the enhancement of both $J_{super}$ and $T_{C}$. The Cr-Te-Cr bond angles are 89.8 and 91.2 for relaxed and strained ML Cr2Ge2Te6, respectively, indicating no significant difference in the hopping parameter $|V_{pd_{1}}|^{2}$ as shown in Table. 1.
Figure 7: The band gap and $T_{C}$ as a function of tensile strain of (a) germanene/BL Cr2Ge2Te6 and (b) germanene/ML Cr2Si2Te6 heterostructures.
Table 1: $|V_{pd_{1}}|^{2}$ and $E_{p^{\prime}}-E_{d_{2}}$ between $5p$ orbitals of Te and $d_{xz}$, $d_{yz}$ orbitals of Cr for relaxed and strained ML Cr2Ge2Te6 (in units of eV).
Strain = 0%
$p_{z}-d_{xz}$ $p_{x}-d_{xz}$ $p_{y}-d_{xz}$ $p_{z}-d_{yz}$ $p_{x}-d_{yz}$ $p_{y}-d_{yz}$
$|V_{pd_{1}}|^{2}$ 0.071 0.012 0.153 0.100 0.222 0.127
$E_{p^{\prime}}-E_{d_{2}}$ 0.621 0.742 0.815 0.628 0.749 0.822
$J_{super}A$ 0.372
Strain = 2%
$p_{z}-d_{xz}$ $p_{x}-d_{xz}$ $p_{y}-d_{xz}$ $p_{z}-d_{yz}$ $p_{x}-d_{yz}$ $p_{y}-d_{yz}$
$|V_{pd_{1}}|^{2}$ 0.057 0.011 0.122 0.086 0.182 0.127
$E_{p^{\prime}}-E_{d_{2}}$ 0.729 0.670 0.743 0.730 0.670 0.743
$J_{super}A$ 0.878
IV QAHE in germanene/BL Cr2Ge2Te6 and germanene/ML Cr2Si2Te6 heterostructures
We also investigated the QAHE in germanene/BL Cr2Ge2Te6 and germanene/ML Cr2Si2Te6 heterostructures, and the results are shown in Fig. 7. Experiments have demonstrated an out-of-plane easy-axis and interlayer ferromagnetic coupling in BL Cr2Ge2Te6 [23]. To determine an appropriate Hubbard $U$ parameter, we calculated the Magnetic anisotropy energy (MAE) $(E_{\parallel}-E_{\perp})$ and the energy difference between AFM and FM interlayer configurations $(E_{AFM}-E_{FM})$ as a function of $U$. As shown in SMs, a value of $U=1.3$ eV accurately reproduces the experimental observations for BL Cr2Ge2Te6. Consequently, this value was adopted for germanene/BL Cr2Ge2Te6 calculations. The SMs also present the four different spin configurations of Cr atoms to calculate the $T_{C}$ of BL Cr2Ge2Te6. Our calculations reveal that the germanene/BL Cr2Ge2Te6 heterostructure exhibits the QAHE up to 64 K under a 1.4% tensile strain as shown in Fig. 7(a).
For the germanene/ML Cr2Si2Te6 heterostructure, a Hubbard $U$ value of 4 eV was employed in our calculations. The similar rescale method is used to counteract the overestimation of $T_{C}$ by MC simulations. Specifically, the calculated $T_{C}^{MC}$(Cr2Si2Te6) of 27 K for ML Cr2Si2Te6 was scaled to the experimentally determined $T_{C}^{exp}$ of 17 K for 2D Cr2Si2Te6 [45]. As shown in Fig. 7(b), the germanene/ML Cr2Si2Te6 heterostructure exhibits the QAHE up to 50 K under a 1.3% tensile strain.
V Conclusion
In this paper, we have investigated the strain-enhanced QAHE in germanene/ML Cr2Ge2Te6, germanene/BL Cr2Ge2Te6 and germanene/ML Cr2Si2Te6 heterostructures based on first principle calculations. Our calculations demonstrate the emergence of topologically nontrivial edge states and quantized anomalous Hall conductance in these heterostructures under tensile strain. We find that increasing tensile strain leads to an increase in $T_{C}$ and a decrease in band gap within these germanene/magnetic semiconductor heterostructures. Consequently, the highest QAHE temperatures achieved are 62 K for the germanene/ML Cr2Ge2Te6 heterostructure under 2.1% tensile strain, 64 K for the germanene/BL Cr2Ge2Te6 heterostructure under 1.4% tensile strain, and 50 K for the germanene/ML Cr2Si2Te6 heterostructure under 1.3% tensile strain. This work identifies experimental available materials for realizing high temperature QAHE states.
Acknowledgements
This work is supported by National Key R&D Program of China (Grant No. 2022YFA1405100), Chinese Academy of Sciences (Grants No. YSBR-030, No. JZHKYPT-2021-08).